# Early-stopping for Transformer model training

### v1

@@ Abstract
This work introduces a novel theoretical framework grounded in Random Matrix Theory (RMT) for analyzing Transformer training dynamics. We focus on the underlying mechanisms that drive performance improvements and derive principled early-stopping criteria. Empirically, we observe that the spectral density of the shallow self-attention matrix V consistently evolves into a heavy-tailed distribution. Utilizing the PL (Power Law) fit to this matrix as a probe, we demarcate training into three stages: structural exploration, heavy-tailed structure stabilization, and convergence saturation. This staging provides guidance for preliminary stopping decisions. Crucially, we propose two consistent and validation-free criteria: a quantitative metric for heavy-tailed dynamics and a novel spectral signature indicative of convergence. The strong alignment between these criteria highlights the utility of RMT for monitoring and diagnosing the progression of Transformer model training.
@@ 1 Introduction
Behind the remarkable success of large language models (LLMs), exemplified by the Transformer architecture, lie substantial training costs and persistent risks of overfitting (Fournier et al., 2023; Variš and Bojar, 2021). According to the well-established scaling law, the performance of an LLM is closely tied to its number of parameters, the quantity of training data, and the computational budget-factors that collectively drive the trend toward ever-larger models (Kaplan et al., 2020). Within this context, early stopping serves as a crucial regularization technique for conserving computational resources and mitigating overfitting once the validation performance saturates(Küken et al., 2025; Wu et al., 2025; Favero et al., 2025).
@@
The classical early stopping approach relies on validation set performance (Prechelt, 2002; Muhammad et al., 2022). However, this strategy exhibits clear limitations, particularly when the training and validation distributions diverge (Taori et al., 2020) or when high-quality validation data are unavailable (Yuan et al., 2025). Heckel and Yilmaz (2020) further revealed the double descent phenomenon, suggesting that naive validation monitoring may fail to capture the true optimal stopping point. Consequently, recent research on early stopping methods has increasingly explored alternative approaches aimed at reducing or even eliminating the reliance on validation sets(Laaouach, 2025; Miseta et al., 2024; Ameryan and Schomaker, 2023). One promising direction is to focus on intrinsic model signals that more accurately reflect a model’s generalization capability. Among these, the spectral distribution of weight matrices-analyzed through the lens of Random Matrix Theory (RMT) has emerged as a powerful analytical tool.
@@
Integrating RMT with regularization techniques enables model compression through spectral pruning without sacrificing accuracy, thereby demonstrating RMT’s potential to enhance deep learning performance (Berlyand et al., 2023, 2025). This principle of separating signal from noise provides a theoretical foundation for understanding and optimizing neural architectures. For many state-of-the-art deep networks, the spectra of weight matrices exhibit heavy-tailed (HT) behavior, well-described by a power-law distribution (Meng and Yao, 2023; Martin and Mahoney, 2021). This phenomenon is not coincidental but reflects an implicit regularization mechanism that emerges during training—known as the Heavy-Tailed Self-Regularization (HT-SR) theory (Martin and Mahoney, 2021). A key corollary of this theory is that “shape metrics” derived from spectral properties can directly predict a model’s generalization ability without access to training or test data. Subsequent studies have confirmed the effectiveness of these metrics for model selection (Yang et al., 2023) and linked them to complex learning behaviors such as “Grokking”, underscoring the potential of spectral analysis (Prakash and Martin, 2025).
@@
From a dynamic viewpoint, HT-SR theory characterizes deep neural network (DNN) training as an evolving process of spectral shape transformation. Martin and Mahoney (2021) proposed a “5+1 phase” framework to describe the progression of implicit self-regularization from weak to strong, while Meng and Yao (2023) later distilled this into a more practical three-stage model, revealing its close relationship to task difficulty. This insight motivates a new type of early stopping criterion: DNN training can be terminated by monitoring changes in the spectral distribution shape.
@@
Despite the Transformer’s widespread adoption (Vaswani et al., 2017), most existing spectral studies have analyzed only the overall weight matrices of Transformer layers (Yang et al., 2023; Martin et al., 2021). Few have investigated the internal spectral dynamics of the Transformer’s core component the attention mechanism specifically its Query (Q), Key (K), and Value (V) matrices. Current Transformer-oriented early stopping methods either continue to rely on validation-based monitoring (Muhammad et al., 2022) or focus on dynamic early-exit mechanisms for accelerating inference (Küken et al., 2025).
The Q, K, and V matrices play pivotal roles in information processing and representation learning within Transformer architectures. Their spectral dynamics are inherently linked to the model’s learning trajectory and ultimate generalization performance. Building upon this observation, we propose a new research direction for early stopping: leveraging the spectral distributions and heavy-tailed characteristics of Q, K, and V matrices to infer the training state and derive a novel early stopping criterion.
@@
    Based on the dynamics of this V matrix, we conceptualize Transformer training as a novel three-phase process: (i) Structural Exploration and Drastic Adjustment, (ii) Heavy-Tailed Structure Formation and Stabilization, and (iii) Performance Saturation and Convergence Plateau.
@@ 2 Formulation
Following the approach of Song et al. (2024), we first theoretically demonstrate that even when $\mathbf{Q}$ and $\mathbf{K}$ are fixed during training and only $\mathbf{V}$ is updated via gradient descent, the cross-entropy loss still converges linearly to the global optimum. This finding supports the hypothesis that the $\mathbf{V}$ matrix plays a more central role in the model than $\mathbf{Q}$ and $\mathbf{K}$.
Since the power law (PL) effectively characterizes the spectral tails of weight matrices exhibiting heavy-tailed behavior (Meng and Yao, 2023; Martin and Mahoney, 2021), we perform PL fitting on all $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices to evaluate the stability and consistency of their heavy-tailed evolution. The results indicate that the $\mathbf{V}$ matrix of the first self-attention layer (en.0.s.a.V) is the most representative, providing empirical validation for our hypothesis.
@@
$\max_{\text{epoch}}{d^{*}-\tilde{d}}$,
@@ 3.1 Models setting and Datasets destails
To investigate the evolution of the Transformer’s core matrices during training, we construct three Transformer variants with distinct architectural configurations, denoted as T1, T2, and T3 models.
  • T1 Model: Consists of three encoder and decoder layers and employs a relatively narrow feed-forward network (d_ff = 800). The total number of parameters is 9.60M.
  • T2 Model: Includes two encoder and decoder layers but widens the feed-forward network to (d_ff = 1800). The total parameter count is 9.62M, which is approximately equal to that of the T1 Model.
  • T3 Model: Comprises six encoder and decoder layers and adopts the same feed-forward dimension as T1 (d_ff = 800), resulting in a total parameter count of 19.20M.
Except for these structural variations, all other hyperparameters remain identical across the models. The complete configuration details are provided in Table 1.
@@
(Param Count) (Encoder & Decoder Layers) (Embedding Dim) (Feed-Forward Dim) (Multi-Head Attn)
@@
To investigate the properties of the weight matrices in the Transformer model, we conduct experiments on the English-to-Chinese (En-Ch) machine translation task. The dataset statistics and relevant details are summarized in Table 2. The translation corpus is obtained from a publicly available parallel dataset [^1] The English–Chinese dataset (CWMT) is available on Baidu AI Studio: https://aistudio.baidu.com/datasetdetail/161162 . In the subsequent experiments, the data are divided into training and validation sets with a 9:1 ratio.
@@
Translation Task Language Tokens (Millions) Vocabulary Size
En-Ch (CWMT) Chinese 11.1 16,011
@@
Consequently, we evaluate our approach using the following three experimental setups.
  • T1_En-Ch9.60M: Use the T1 model trained on the En-Ch dataset.
  • T2_En-Ch9.62M: Use the T2 model trained on the En-Ch dataset.
  • T3_En-Ch19.20M: Use the T3 model trained on the En-Ch dataset.
@@ 3.2 Selection of Core Matrix by Power Law distribution for Transformer
According to Random Matrix Theory (RMT), the spectral distribution of an untrained random matrix theoretically follows the Marchenko–Pastur (MP) distribution. However, our experiments reveal that even in the early stages of training, the weight matrices of Transformer models deviate substantially from the MP distribution. Instead, their empirical spectral densities (ESDs) exhibit prominent heavy-tailed behavior. This observation suggests that the MP distribution is not an adequate descriptor of weight-matrix properties in large language models (Prakash and Martin, 2025). To better capture this intrinsic heavy-tailed nature, we fit the ESDs of key matrices using a power-law distribution and use this fit as the criterion for identifying representative matrices. Specifically, this study evaluates the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices along two dimensions:
  • Significance of heavy-tailed characteristics: Identify the matrix that exhibits the strongest heavy-tailed signature via spectral analysis, thereby justifying the use of the power-law model for quantitative characterization.
  • Stability of dynamic evolution: Select the matrix whose power-law parameters evolve most smoothly and consistently during training.
(Gurbuzbalaban et al., 2021) established that the heavy-tailed phenomenon observed in the eigenvalue spectra of deep neural network weight matrices reflects an intrinsic regularization effect of gradient-descent-based optimization. (Song et al., 2024) further elucidated the distinct optimization dynamics among different weight matrices in Transformer models under quadratic loss. Building upon these theoretical foundations, we propose Research Proposition A.1 (see Appendix A), demonstrating that analogous conclusions extend to machine translation tasks. Specifically, in a single-layer Transformer under over-parameterization (i.e., sufficiently large embedding dimension D) and proper initialization, when $\mathbf{Q}$ and $\mathbf{K}$ remain fixed while only $\mathbf{V}$ undergoes gradient descent updates, the cross-entropy loss converges linearly to the global optimum. This behavior originates from the softmax kernel in the attention mechanism: while the optimization trajectories of $\mathbf{Q}$ and $\mathbf{K}$ are nonlinearly coupled with the kernel, the dynamics of $\mathbf{V}$ remain approximately linear, thus providing a cleaner observational window.
@@
Following Alstott et al. (2014), we posit that if a power-law model does not significantly outperform an exponential model in fitting empirical data, there is insufficient evidence to classify the distribution as heavy-tailed. Consequently, we employ the exponential distribution as a competitive baseline and use log-likelihood ratio testing to determine which model more accurately captures heavy-tailed phenomena in the spectral data.
For a given attention matrix $\mathbf{W}$, the testing procedure is implemented as follows:
@@
For the continuous power-law probability density function (PDF),
@@
Step 3: Model Comparison
The difference between the log-likelihoods of the two models, denoted as the log-likelihood ratio $\mathcal{R}$, is computed as:
$\mathcal{R}=\ln\frac{\mathcal{L}_{1}}{\mathcal{L}_{2}}=\sum_{i=1}^{n}[\ln p_{1}(x_{i})-\ln p_{2}(x_{i})].$ (3.3)
If $\mathcal{R}>0$, the data are more likely to follow a power-law distribution than an exponential one, providing evidence of heavy-tailed behavior. Otherwise, heavy-tailedness cannot be inferred.
To assess the statistical significance of this result, we adopt the procedure proposed by Clauset et al. (2009) to compute a $p$-value. If $p<0.1$ (a commonly used significance threshold), the empirical data are considered significantly more consistent with a power-law distribution, indicating that the matrix exhibits statistically significant heavy-tailed spectral characteristics. Due to space limitations, detailed fitting results are presented in Appendix B.
We applied this procedure to the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices across all attention layers of the Transformer models (T1_En-Ch9.60M, T2_En-Ch9.62M). Results concerning the stability of dynamic evolution are provided in Appendix C.
@@ 3.3 Three Training Phases for Transformer model training
The analysis in Section B of the Appendix has shown that the fitted parameters of the model begin to stabilize after approximately 50 epochs. To capture this transition precisely—from an initial fluctuating phase to a convergent state—we analyze the first 205 training epochs of this layer, focusing on the evolution of its PL fitting parameters. The detailed trajectories of the fitted parameters ($\alpha$ and $x_{min}$) across multiple epochs are presented in Figure 1.
Figure 1: Evolution of $\alpha$ and $x_{\text{min}}$ for en.0.s.a.v matrix. Figure 1 illustrates the evolution of the PL fitting parameters for the en.0.s.a.v matrix, tracked across the three experimental configurations (T1_En-Ch9.60M, T2_En-Ch9.62M, T3_En-Ch19.20M) as a function of training epochs. The blue solid line (left Y-axis) represents the parameter $\alpha$, while the red dashed line (right Y-axis) corresponds to the PL lower bound parameter $x_{\text{min}}$.
@@
  • Phase I: Structural Exploration and Drastic Adjustment. During the initial training epochs, the spectral parameters undergo the most substantial dynamic changes. Both the PL exponent $\alpha$ and the minimum eigenvalue threshold $x_{\min}$ remain elevated and exhibit considerable volatility. This indicates that the model is transitioning from its random initialization, with weight matrices undergoing large-scale, global structural reorganization to rapidly explore an effective feature space.
  • Phase II: Heavy-Tailed Structure Formation and Stabilization. Following the initial exploration, the system enters a prolonged phase of stable evolution. The hallmark of this phase is the convergence of $\alpha$ to a low, stable value (consistently below 3 and approximately 2.5 across all experiments), maintained with high stability over hundreds of epochs. This stabilization indicates the formation of a stable heavy-tailed spectral structure within the weight matrices, a key mechanism for achieving Implicit Regularization in deep learning models. Concurrently, $x_{\min}$ decreases and stabilizes at a low level (around 0.1).
  • Phase III: Performance Saturation and Convergence Plateau. In the later training stages, the model enters a performance saturation phase. The most notable parametric change is a gradual upward trend in $x_{\min}$, coinciding with the increasing difficulty of achieving additional performance gains. This suggests that the primary feature-learning process has largely completed, and subsequent training primarily fine-tunes the model, indicating a convergence plateau.
Based on the evolutionary patterns illustrated in Figure 1, the training processes of the three experimental configurations can be quantitatively segmented into distinct phases. The corresponding phase durations and characteristics are summarized in the Table 3 below.
Table 3: Summary of Training Phase Divisions for Each Experimental Configuration.
Experimental
Configuration
Phase I:
Structural Exploration
and Drastic Adjustment
Phase II:
Heavy-Tailed Structure
Formation and Stabilization
Phase III:
Performance Saturation
and Convergence Plateau
T1_En-Ch 9.60M Epochs 1–25 Epochs 26–105 Epochs 150–205
T2_En-Ch9.62M Epochs 1–27 Epochs 28–155 Epochs 200–205
T3_En-Ch19.20M Epochs 1–24 Epochs 25–55 Epochs 100–205
@@
To further validate the correspondence between these internal spectral dynamics and external learning performance, Figure 3 shows the evolution of Loss and Accuracy values across all training epochs for the three experimental setups. During Phase I, the model exhibits the most rapid performance improvement, characterized by a sharp decline in loss and a rapid increase in accuracy. In Phase II, the rate of improvement slows, and the curves become smoother as the model undergoes steady learning and fine-tuning. Finally, in Phase III, the performance curves plateau, with both validation loss and accuracy showing negligible further improvement, consistent with the microscopic phenomena observed in the spectral tail.
Therefore, for Transformer models, the optimal early-stopping window lies between the late portion of Phase II and the early portion of Phase III—after the heavy-tailed structure has stabilized but before the model reaches full performance saturation.
3.4 Recognition criteria for Heavy-Tail based on PL fitting
This section presents a novel spectral criterion based on PL fitting to more accurately identify the phenomenon of Heavy-Tailed period in the en.0.s.a.v matrix during Transformer model training. It is extremely important for us to provide early stopping criteria to guide model training and manage model performance expectations in the future. We focus on tail characteristics of en.0.s.a.v, specifically, the portion where the eigenvalues $x\geq x_{\min}$. All $n_{\text{tail}}$ data points satisfying this condition can be regarded as being independently and identically drawn (i.i.d.) from an ideal PL distribution.
Based on Eq (3.1), the Cumulative Distribution Function (CDF) of this idealized power-law distribution is defined as follows,

### v2

@@ Abstract
This work, based on Random Matrix Theory (RMT), introduces a novel early-stopping strategy for Transformer training dynamics. Utilizing the Power Law (PL) fit to tansformer attention matrices as a probe, we demarcate training into three stages: structural exploration, heavy-tailed structure stabilization, and convergence saturation. Empirically, we observe that the spectral density of the shallow self-attention matrix $V$ consistently evolves into a heavy-tailed distribution. Crucially, we propose two consistent and validation-set-free criteria: a quantitative metric for heavy-tailed dynamics and a novel spectral signature indicative of convergence. The strong alignment between these criteria highlights the utility of RMT for monitoring and diagnosing the progression of Transformer model training.
@@ 1 Introduction
Behind the remarkable success of large language models (LLMs), exemplified by the Transformer architecture, lie substantial training costs and persistent risks of overfitting (fournier2023practical; varis2021sequence). According to the well-established scaling law, the performance of an LLM is closely tied to its number of parameters, the quantity of training data, and the computational budget-factors that collectively drive the trend toward ever-larger models (kaplan2020scaling). Within this context, early stopping serves as a crucial regularization technique for conserving computational resources and mitigating overfitting once the validation performance saturates(kuken2025early; wu2025benefits; favero2025bigger).
@@
The classical early stopping approach relies on validation set performance (prechelt2002early; muhammad2022early). However, this strategy exhibits clear limitations, particularly when the training and validation distributions diverge (taori2020measuring) or when high-quality validation data are unavailable (yuan2024early). heckel2021early further revealed the double descent phenomenon, suggesting that naive validation monitoring may fail to capture the true optimal stopping point. Consequently, recent research on early stopping methods has increasingly explored alternative approaches aimed at reducing or even eliminating the reliance on validation sets (laaouach2025halt; miseta2024surpassing; ameryan2023limit). One promising direction is to focus on intrinsic model signals that more accurately reflect a model’s generalization capability. Among these, the spectral distribution of weight matrices-analyzed through the lens of Random Matrix Theory (RMT) has emerged as a powerful analytical tool.
@@
Integrating RMT with regularization techniques enables model compression through spectral pruning without sacrificing accuracy, thereby demonstrating RMT’s potential to enhance deep learning performance (berlyand2024enhancing; berlyand2025pruning). This principle of separating signal from noise provides a theoretical foundation for understanding and optimizing neural architectures. For many state-of-the-art deep networks, the spectra of weight matrices exhibit heavy-tailed (HT) behavior, well-described by a power-law distribution (meng2023impact; martin2021implicit). This phenomenon is not coincidental but reflects an implicit regularization mechanism that emerges during training—known as the Heavy-Tailed Self-Regularization (HT-SR) theory (martin2021implicit). A key corollary of this theory is that “shape metrics” derived from spectral properties can directly predict a model’s generalization ability without access to training or test data. Subsequent studies have confirmed the effectiveness of these metrics for model selection (yang2023test) and linked them to complex learning behaviors such as “Grokking”, underscoring the potential of spectral analysis (prakash2025grokking).
@@
From a dynamic viewpoint, HT-SR theory characterizes deep neural network (DNN) training as an evolving process of spectral shape transformation. martin2021implicit proposed a “5+1 phase” framework to describe the progression of implicit self-regularization from weak to strong, while meng2023impact later distilled this into a more practical three-stage model, revealing its close relationship to task difficulty. This insight motivates a new type of early stopping criterion: DNN training can be terminated by monitoring changes in the spectral distribution shape.
@@
Despite the Transformer’s widespread adoption (vaswani2017attention), most existing spectral studies have analyzed only the overall weight matrices of Transformer layers (yang2023test; martin2021predicting). Few have investigated the internal spectral dynamics of the Transformer’s core component the attention mechanism specifically its Query (Q), Key (K), and Value (V) matrices. Current Transformer-oriented early stopping methods either continue to rely on validation-based monitoring (muhammad2022early) or focus on dynamic early-exit mechanisms for accelerating inference (kuken2025early).The Q, K, and V matrices play pivotal roles in information processing and representation learning within Transformer architectures (anonymous2025from; borji2023key). Their spectral dynamics are inherently linked to the model’s learning trajectory and ultimate generalization performance. Building upon this observation, we propose a new research direction for early stopping: leveraging the spectral distributions and heavy-tailed characteristics of Q, K, and V matrices to infer the training state and derive a novel early stopping criterion.
We summarize our contributions as follows:
@@
    Based on the dynamics of en.0.s.a.V matrix, we conceptualize Transformer training as a novel three-phase process: (i) Structural Exploration and Drastic Adjustment, (ii) Heavy-Tailed Structure Formation and Stabilization, and (iii) Performance Saturation and Convergence Plateau.
@@ 2 Formulation
Since the power law (PL) effectively characterizes the spectral tails of weight matrices exhibiting heavy-tailed behavior (meng2023impact; martin2021implicit), we perform PL fitting on all $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices to evaluate the stability and consistency of their heavy-tailed evolution. The results indicate that the $\mathbf{V}$ matrix of the first self-attention layer (en.0.s.a.V) is the most representative, providing empirical validation for our hypothesis.
@@
$\max_{\text{epoch}}{\{d^{*}-\tilde{d}\}}$,
@@ 3.1 Models setting and Datasets destails
To investigate the evolution of the Transformer’s core matrices during training, we construct three Transformer variants with distinct architectural configurations, denoted as T1, T2, and T3 models. Except for these structural variations, all other hyperparameters remain identical across the models. The complete configuration details are provided in Table 1.
@@
To investigate the properties of the weight matrices in the Transformer model, we conduct experiments on the English-to-Chinese (En-Ch) machine translation task. The dataset statistics and relevant details are summarized in Table 2. The translation corpus is obtained from a publicly available parallel dataset. [^1] The English–Chinese dataset (CWMT) is available on Baidu AI Studio: https://aistudio.baidu.com/datasetdetail/161162 In the subsequent experiments, the data are divided into training and validation sets with a 9:1 ratio.
@@
Task Language Tokens (Millions) Vocab Size
En-Ch Chinese 11.1 16,011
@@ 3.2 Selection of Core Matrix by Power Law distribution for Transformer
According to Random Matrix Theory (RMT), the spectral distribution of an untrained random matrix theoretically follows the Marchenko–Pastur (MP) distribution. However, our experiments reveal that even in the early stages of training, the weight matrices of Transformer models deviate substantially from the MP distribution. Instead, their empirical spectral densities (ESDs) exhibit prominent heavy-tailed behavior. This observation suggests that the MP distribution is not an adequate descriptor of weight-matrix properties in large language models (prakash2025grokking). To better capture this intrinsic heavy-tailed nature, we fit the ESDs of key matrices using a power-law distribution and use this fit as the criterion for identifying representative matrices. Specifically, this study evaluates the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices along two dimensions: Significance of heavy-tailed characteristics: Identify the matrix that exhibits the strongest heavy-tailed signature via spectral analysis, thereby justifying the use of the power-law model for quantitative characterization. Stability of dynamic evolution: Select the matrix whose power-law parameters evolve most smoothly and consistently during training.
gurbuzbalaban2021heavy established that the heavy-tailed phenomenon observed in the eigenvalue spectra of deep neural network weight matrices reflect an intrinsic regularization effect of gradient-descent-based optimization. song2024unraveling further elucidated the distinct optimization dynamics among different weight matrices in Transformer models under quadratic loss. Building upon these theoretical foundations, we propose Proposition A.1 (see Appendix A), demonstrating that analogous conclusions extend to machine translation tasks. Specifically, in a single-layer Transformer under over-parameterization (i.e., sufficiently large embedding dimension D) and proper initialization, when $\mathbf{Q}$ and $\mathbf{K}$ remain fixed while only $\mathbf{V}$ undergoes gradient descent updates, the cross-entropy loss converges linearly to the global optimum. This behavior originates from the softmax kernel in the attention mechanism: while the optimization trajectories of $\mathbf{Q}$ and $\mathbf{K}$ are nonlinearly coupled with the kernel, the dynamics of $\mathbf{V}$ remain approximately linear, thus providing a cleaner observational window.
@@
For a given attention matrix $\mathbf{W}$, the fitting procedure is implemented as follows:
@@
Two candidate distributions are fitted to this eigenvalue tail. Using maximum likelihood estimation (MLE), we estimate the optimal exponent $\alpha$ for the power-law model and the decay rate $\lambda$ for the exponential model, respectively.
For the continuous power-law probability density function,
@@
For the continuous exponential probability density function,
$p(x;\lambda,x_{\min})=\lambda e^{-\lambda(x-x_{\min})},$ (3.3)
where $\lambda$ represents the decay rate, and $x_{\min}$ denotes the lower threshold as defined above. The MLE for $\lambda$ is given by:
$\hat{\lambda}=n\left[\sum_{i=1}^{n}(x_{i}-x_{\min})\right]^{-1}.$ (3.4)
Following alstott2014powerlaw, we posit that if a power-law model does not significantly outperform an exponential model in fitting empirical data, there is insufficient evidence to classify the distribution as heavy-tailed. Consequently, we employ the exponential distribution as a competitive baseline and use log-likelihood ratio testing to determine which model more accurately captures heavy-tailed phenomena in the spectral data.If $\mathcal{R}>0$, the data are more likely to follow a power-law distribution than an exponential one, providing evidence of heavy-tailed behavior. Otherwise, heavy-tailedness cannot be inferred.To assess the statistical significance of this result, we adopt the procedure proposed by clauset2009power to compute a $p$-value. If $p<0.1$ (a commonly used significance threshold), the empirical data are considered significantly more consistent with a power-law distribution, indicating that the matrix exhibits statistically significant heavy-tailed spectral characteristics.
We summarize the procedure in Algorithm 1. Due to space constraints, detailed fitting results are deferred to Appendix B.1. Crucially, these results reveal that the en.0.s.a.v matrix consistently exhibits pronounced heavy-tailed characteristics throughout the training process. This empirical evidence fulfills our objective of identifying the matrix with the strongest heavy-tailed signature.
Algorithm 1 Test for Power-Law Superiority
Input: Weight matrix $\mathbf{W}$, Significance threshold $p_{\text{thres}}=0.1$
Output: is_heavy_tailed, $x_{\min}$, $\hat{\alpha}$, $\hat{\lambda}$, $\mathcal{R}$, $p\text{-value}$
Step 1: Eigenvalue Selection
  $\mathbf{X}=\mathbf{W}^{\top}\mathbf{W}$
  $\text{Eigval}=\text{compute\_eigval}(\mathbf{X})$
  $x_{\min}=\text{ks\_test}(\text{Eigval},\text{target=``power\_law''})$
  $\text{Eigval\_tail}=\{\xi\in\text{Eigval}\mid\xi\geq x_{\min}\}$
Step 2: Model Fitting
  $\hat{\alpha}=\text{mle\_pl}(\text{Eigval\_tail},x_{\min})$
  $\hat{\lambda}=\text{mle\_exp}(\text{Eigval\_tail})$
Step 3: Model Comparison and Heavy-Tailedness Test
  $\log\mathcal{L}_{\text{PL}}=\sum_{\xi\in\text{Eigval\_tail}}\ln\left(\text{PL}_{\text{PDF}}(\xi,\hat{\alpha},x_{\min})\right)$
[…]
  $\mathcal{R}=\ln\mathcal{L}_{\text{PL}}-\ln\mathcal{L}_{\text{Exp}}$
[…]
  $\text{is\_heavy\_tailed}=(\mathcal{R}>0)\land(p\text{-value}<p_{\text{thres}})$
Return $\text{is\_heavy\_tailed},x_{\min},\hat{\alpha},\hat{\lambda},\mathcal{R},p\text{-value}$
Furthermore, we performed power-law parameter estimation on the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices across all attention layers of the Transformer models (T1_En-Ch9.60M, T2_En-Ch9.62M). The corresponding results regarding the stability of dynamic evolution are detailed in Appendix C.
@@ 3.3 Three Training Phases for Transformer model training
The analysis in Appendix B.1 has shown that the fitted parameters of the model begin to stabilize after approximately 50 epochs. To capture this transition precisely—from an initial fluctuating phase to a convergent state—we analyze 1-55,100-105,150-155,200-205 training epochs of this layer, focusing on the evolution of its PL fitting parameters. The detailed trajectories of the fitted parameters ($\alpha$ and $x_{min}$) for the en.0.s.a.v matrix, tracked under the experimental configuration T3_En-Ch19.20M as a function of training epochs, are shown in Figure 1; the corresponding plots for T1_En-Ch9.60M and T2_En-Ch9.62M [^3] T1_En-Ch9.60M: T1 model (En-Ch dataset). T2_En-Ch9.62M: T2 model (En-Ch dataset). T3_En-Ch19.20M: T3 model (En-Ch dataset). are provided in Figure 7 from Appendix B.2.
Figure 1: Evolution of $\alpha$ and $x_{\text{min}}$ for en.0.s.a.v matrix of T2_En-Ch9.62M. The blue solid line (left Y-axis) represents the parameter $\alpha$, while the red dashed line (right Y-axis) corresponds to the PL lower bound parameter $x_{\text{min}}$.
The power-law (PL) fitting parameter $\alpha$ is a key metric for evaluating model training stability and feature learning quality martin2021predicting. To investigate the consistency of model training evolution under different parameter settings, we conducted a density statistical analysis on the $\alpha$ values derived from PL fitting of the en.0.s.a.v matrices across three groups of experiments (T1, T2, T3), with the corresponding density histograms shown in Figure 2(a). The results indicate that the $\alpha$ values of the three experiments are mainly concentrated in the range of 2–4. Among them, the density histograms of Experiments T1 and T2 (with identical parameter settings) exhibit a high degree of overall similarity, demonstrating that under consistent parameter configurations, the model’s training evolution process possesses extremely high reproducibility and similarity. In contrast, during the training of Model T3 (with different parameter settings), obvious outliers of $\alpha$ values in the range of 10–13 were observed, these outliers correspond to the initial stage of model training.The T3 model with a larger parameter scale requires significant global structural reorganization to complete the initial construction of an effective feature space, a process that directly leads to the abnormal deviation of $\alpha$ values.
The optimal fitting threshold $x_{min}$ in the power-law (PL) fitting process varies with the dynamic update of the en.0.s.a.v matrix, while in the later stages of model training, $x_{min}$ exhibits reduced fluctuations and tends to stabilize. To eliminate the interference of $x_{min}$ fluctuations on the statistical results of $\alpha$ and ensure the consistency of the analysis, we fixed $x_{min}=0.1$ and re-performed PL fitting and density statistics. At this point, the PL exponents ($\alpha$) of the en.0.s.a.v matrices for the three models are mainly concentrated between 1.8 and 2.6, showing a more centralized distribution, with the corresponding density distribution diagrams presented in Figure 2(b).
(a) No fix $x_{min}$
(b) Fix $x_{min}=0.1$
Figure 2: PL epxonents density histograms of T1, T2, and T3 models
From the statistical results after fixing $x_{min}=0.1$ (selected based on the approximate empirical values from the stable heavy-tailed phase of the three experimental groups), the $\alpha$ distributions of Models T1 and T2 (with the same parameter settings) highly overlap, indicating that the feature extraction logic and effects of models with identical parameters have good consistency. In contrast, the $\alpha$ values of Model T3 are generally larger, showing a significant right-shift feature in its distribution, but the distribution patterns of the three models still maintain a certain degree of similarity. The above results confirm that although there are slight differences in feature extraction accuracy and the extent of generalization ability improvement among models with different parameters, the evolutionary laws of the core feature extraction process have inherent consistency. This empirical result provides critical support for the rationality and generality of the following three stages of model training and the unified early stopping guidance framework proposed in this paper.
@@
Phase I: Structural Exploration and Drastic Adjustment.
During the initial training epochs, the spectral parameters undergo the most substantial fluctuations. Both the PL exponent $\alpha$ and the minimum eigenvalue threshold $x_{\min}$ remain elevated and exhibit considerable volatility. This behavior indicates that the model is transitioning away from its random initialization, undergoing large-scale global structural reorganization to rapidly explore and locate an effective feature space.
Phase II: Heavy-Tailed Structure Formation and Stabilization. Following the initial exploration, the system enters a prolonged phase of stable evolution. The hallmark of this phase is the convergence of $\alpha$ to a low, stable value (consistently $\alpha<3$, clustering around $2.5$), which is maintained over hundreds of epochs. This stabilization signals the formation of a robust heavy-tailed spectral structure within the weight matrices—a key mechanism linked to Implicit Regularization in deep learning. Concurrently, $x_{\min}$ decreases and stabilizes at a low magnitude (approximately $0.1$).
Phase III: Performance Saturation and Convergence Plateau. In the later training stages, the model enters a saturation phase. The most notable parametric signature is a gradual upward trend in $x_{\min}$, which correlates with the diminishing returns in performance gains. This suggests that the primary feature-learning process is largely complete; subsequent training serves primarily to fine-tune the model, indicating a convergence plateau.
Based on the evolutionary patterns illustrated in Figure 1, the training processes for the three experimental configurations can be quantitatively segmented into these distinct phases. The corresponding phase durations and characteristics are summarized in Table 3
Table 3: Epoch-Based Division of Training Stages for Each Model
Model
Phase I
Phase II
Phase III
T1 1–25 26–105 150–205
T2 1–27 28–155 200–205
T3 1–24 25–55 100–205
@@
In Appendix B.2, we compare the three divided training phases with the variations of loss values and accuracy during the model training process, thereby further verifying the effectiveness of our phase division criteria.
Through the above analysis, we can conclude that, for Transformer models, the optimal early-stopping window lies between the late portion of Phase II and the early portion of Phase III—after the heavy-tailed structure has stabilized but before the model reaches full performance saturation.
3.4 Construction of a Power-Law-Based Heavy-Tailed Metric: Key Theoretical Details
This section formulates a quantitative spectral metric based on the goodness-of-fit of the Power-Law distribution, establishing a statistical threshold to rigorously distinguish heavy-tailed characteristics.
We focus on tail characteristics of en.0.s.a.v, specifically, the portion where the eigenvalues $x\geq x_{\min}$. All $n_{\text{tail}}$ data points satisfying this condition can be regarded as being independently and identically drawn (i.i.d.) from an ideal PL distribution. Based on Eq (3.1), the Cumulative Distribution Function (CDF) of this idealized power-law distribution is defined as follows,